OpenStudy (anonymous):

Kinda confused on this one and others like it . Show how to solve step by step? Ryan flew from Wiley Post to Ponca City and back. Ryan maintained an average rate of 450 mph going to Ponca City and an average rate of 400 mph returning to Wiley Post. If the actual flying time for the roundtrip was 1 hour, about how far is it from Wiley Post to Ponca City? Round to nearest mile.

OpenStudy (anonymous):

Distance = time x speed

OpenStudy (agreene):

\[\frac{miles}{hour}*hour \rightarrow\frac{miles}{\cancel{hour}}*\cancel{hour}=miles\] So, (450+400)mph*1h = 850miles

OpenStudy (anonymous):

Let the time for the first trip be x so the time of returning is 1-x 450x = 400 (1-x) 450x = 400 - 400x 950x = 400 x = 950/400 450 x (950/400) = 1068.75 miles

OpenStudy (anonymous):

@agreene for different speed they have different time

OpenStudy (asnaseer):

@moneybird - I think you made a small error in your calculation for x. from the line:\[950x=400\]the next line should be:\[x=400/950\]

OpenStudy (anonymous):

yup my mistake. x = 400/950 450 x (400/950) = 3600/19

OpenStudy (asnaseer):

@moneybird - sorry - you also made one other error: 400+450=850 (not 950)

OpenStudy (anonymous):

o yeah thanks for pointing that out calculation mistake haha 450 x (400/850) = 3600/17

OpenStudy (asnaseer):

For the benefit of @Cj Sade, I will try and summarise this:  t1 is the time taken to fly from Wiley Post to Ponca City at 450 mph t2 is the time taken to fly from Ponca City to Wiley Post at 400 mph we know the total time of the flight is 1 hour so t1 + t2 = 1, giving: (1) t2 = 1 - t1 now, distance = speed * time taken, so, when flying from Wiley to Ponca, the distance flown is: (2) d = 450 * t1 and, when flying from Ponca to Wiley, the distance flown is the same, so: (3) d = 400 * t2 = 400 * (1 - t1) from equation (1) above we now have two equations for d - equations (2) and (3), so we can equate them to get: 450 * t1 = 400 * (1 - t1) = 400 - (400 * t1), therefore: 850 * t1 = 400 (4) t1 = 400 / 850 = 40 / 85 = 8 / 17 so finally, we can use equation (2) to work out the distance as follows: d = 450 * t1 = 450 * 8 / 17 = 3600 / 17 = 211.77 miles (approx)